The driver of a car travels 10 km, north after that he turned 8 km, east to go to his work. How far he might travel if he take the short cut straight to his work? Express your answer in vector.
1
Expert's answer
2020-09-14T10:19:46-0400

If we direct the y-axis along the northern direction and x-axis along the eastern one, then the coordinates of the first part vector will be:


"\mathbf{d_1} = (0,10) \space [km]"

The coordinates of the second part vector will be:


"\mathbf{d_2} = (8,0) \space[km]"

The short path vector will be:


"\mathbf{d} = \mathbf{d_1} + \mathbf{d_2} = (0,10) + (8,0) = (8,10) \space[km]"

The lenght of the short path is:


"d = |\mathbf{d}| = \sqrt{8^2 + 10^2} \approx 12.8\space km"

Answer. Vector of the short path: "\mathbf{d} = (8,10) \space[km]". Length of the short path: "d = 12.8 \space km".